I'm having a hard time with this question and I just help to know how to solve it.
What is 3 times the y-coordinate of the solution to the following system? { -x+y=8
y=4x

Answer by jim_thompson5910(35256)

First find the solution to the system (x,y)

-x+y=8

-x+4x=8 ... replace y with 4x. This works because y = 4x

3x = 8

x = 8/3

---------------

Plug x = 8/3 into y = 4x to find y


y = 4x

y = 4(8/3)

y = 32/3


------------------------

Now multiply y = 32/3 by 3 to get "3 times the y-coordinate"

3y = 3*(32/3) = 96/3 = 32

=================================================================

The final answer is 32